Given a mean (8.0 hours) and a standard deviation (4.5 hours) of a population, calculate the probability of an event that falls between 5 to 10 hours. This question is due on day 5.
Answer by stanbon(75887)

Given a mean (8.0 hours) and a standard deviation (4.5 hours) of a population, calculate the probability of an event that falls between 5 to 10 hours. This question is due on day 5.
-----------------
Find the z-score of 5 and of 10:
z(5) = (5-8)/4.5 = -3/4.5 = -1/1.5 = -0.667
z(10)= (10-8)/4.5 = 2/4.5 = 0.4444
-------------
Use your z-chart or your calculator to find the following:
Probability (5<=x<=10) = P(-0.667 <= z <= 0.444) = 0.4191
===============
Cheers,
Stan H.